OpenStudy (anonymous):

Can anyone help me with solving real world problems for Algebra II ? Please & thank you .

OpenStudy (ghazi):

is that your question?

OpenStudy (anonymous):

lol, no.. this is one of my questions : A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air?

OpenStudy (ghazi):

well d=constant in both the cases from orlando to denever and back again so just equate two distances you'll get speed of airplane

OpenStudy (anonymous):

Hartnn: I believe so .. & Ghazi: ... how would I do that?

OpenStudy (ghazi):

well i am getting speed of airplane = 40 miles/hour is it correct lol..

OpenStudy (ghazi):

@amistre64 help me to figure this out

hartnn (hartnn):

Now since distance covered is equal,we will equate speed* time for both Let S be the speed of airplane in still air,and W=40 be the speed of the wind for 1st journey,total speed S1= S+W ,total time T1=5 for 2nd journey,total speed S2=S-W,total time T2=6 As speed = distance/time and since distances are equal, S1*T1=S2*T2 so (S+40)*5=(S-40)*6 can u solve for S now?? @gabbytheflower

OpenStudy (amistre64):

assuming the same route is taken to and fro; distance remains constant distance = (unitspeed + 40) * 5 distance = (unitspeed - 40) * 6

OpenStudy (amistre64):

80*5 = 0*6 ? nah

OpenStudy (ghazi):

what's this?

OpenStudy (amistre64):

"well i am getting speed of airplane = 40 miles/hour" (40+40)*5 (40-40)*6 just saying :)

OpenStudy (amistre64):

at least theres no yeild signs up there ....

OpenStudy (ghazi):

ohkay :)

OpenStudy (ghazi):

:))